OpenStudy (anonymous):

The function T(n) is defined below. Which of the following are equal to T(10)? Check ALL that apply. T(n)=6n-4 A)T(9)+6 B)64 C)56 D)T(8)+T(2)

OpenStudy (anonymous):

T(10) = 6*(10) - 4 = 60 - 4 = ... there is more than 1 answer... btw...

OpenStudy (anonymous):

one is 56, what is the other one?

OpenStudy (anonymous):

do the computation like i showed...

OpenStudy (anonymous):

I have been working on this for an hour, I am a 45 year old single mother trying to help her 15 year old son. I know one answer is 56, I think the other one is T(8)+T(2). Is that it?